OpenStudy (anonymous):

Find the equation of the line below. Arrange your answer in the form y = mx + b, where b is the constant. (Equation of line passing through point A = ( 2 , 2 ) , and point B = ( −2 , −2 ) is)

OpenStudy (anonymous):

I put them into my graphing calculator, and got this: so, y = x + 0?

OpenStudy (anonymous):

originally, it was -x + y = 0

OpenStudy (anonymous):

I think so. My answer was y=1 but was incorrect.

OpenStudy (anonymous):

hmmm....i don't know what to tell you honestly, ask a person who has 99. Sorry ;_____;

OpenStudy (anonymous):

I tried y = 0 but was also incorrect.

OpenStudy (anonymous):

hmmm...I'm sorry I don't know what to say. I tried.

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

ill take a look hold on

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

i got y=1x+0 because you subtract the ys and xs and get -4/-4 which is 1. so that is your slope. you then put it in the format. y=1x+b if you plug in one coordinate (2,2) for x and y you get 2=2+b so be can only be 0

OpenStudy (anonymous):

so y=x

OpenStudy (anonymous):

i hope that helped

OpenStudy (anonymous):

Thank's for the help

OpenStudy (anonymous):

welcome, if you try the second set of numbers it works as well.